Margaret invested $15,000 in a savings account and an IRA. The savings account yielded 3% while the IRA yielded 5%. If her total revenue was $550, how much did she invest in each?
Answer by addingup(3677)

Savings: 0.03x
IRA: (15000-x)*0.05
Put it all together:
0.03x+(15000-x)*0.05= 550
0.03x+750-0.05x= 550 Subtract 750 on both sides and add/subtract x on left:
-0.02x= -200 Now divide both sides by -0.02 and remember -/- = +
x= 10000 _o 10,000 are invested at 3% and 15000-10000= 5000 at 5%
--------------------------------------------------
Check:
10000*0.03= 300
5000*0.05= 250
300+250= 550 We have the correct answer